Answer all assigned questions and problems, and show all work.

  1. Determine the pH of (a) a 0.40 MCH3CO2H solution, (b) a solution that is 0.40 MCH3CO2H and 0.20 MNaCH3CO2. (8 points)

(Reference: Chang 16.5)

  1. Which of the following solutions can act as a buffer? (a) KCl/HCl, (b) KHSO4/H2SO4, (c) KNO2/HNO2.(5 points)

(Reference: Chang 16.9)

  1. Calculate the pH of the buffer system made up of 0.15 MNH3/0.35 MNH4Cl. (5 points)

(Reference: Chang 16.11)

  1. The pH of a bicarbonate-carbonic acid buffer is 8.00. Calculate the ratio of the concentration of carbonic acid (H2CO3) to that of the bicarbonate ion (HCO3). (5 points)

(Reference: Chang 16.13)

  1. The pH of a sodium acetate–acetic acid buffer is 4.50. Calculate the ratio [CH3COO]/[CH3COOH].(5 points)

(Reference: Chang 16.15)

  1. Calculate the pH of the 0.20 MNH3/0.20 MNH4Cl buffer. What is the pH of the buffer after the addition of 10.0 mL of 0.10 MHCl to 65.0 mL of the buffer? (8 points)

(Reference: Chang 16.17)

  1. A 0.2688 g sample of a monoprotic acid neutralized 16.4 mL of 0.08133 M KOH solution. Calculate the molar mass of the acid. (5 points)

(Reference: Chang 16.27)

  1. In a titration experiment, 12.5 mL of 0.500 MH2SO4neutralize 50.0 mL of NaOH. What is the concentration of the NaOH solution? (5 points)

(Reference: Chang 16.29)

  1. A 0.1276 g sample of an unknown monoprotic acid was dissolved in 25.0 mL of water and titrated with 0.0633 MNaOH solution. The volume of base required to bring the solution to the equivalence point was 18.4 mL. (a) Calculate the molar mass of the acid. (b) After 10.0 mL of base had been added during the titration, the pH was determined to be 5.87. What is the Kaof the unknown acid? (10 points)

(Reference: Chang 16.31)

            

  1. Calculate the pH at the equivalence point for the following titration: 0.20 MHCl versus 0.20 Mmethylamine (CH3NH3;Kb= 4.4 × 10–4)(5 points)

            

(Reference: Chang 16.33)

  1. A 25.0 mL solution of 0.100 MCH3COOH is titrated with a 0.200 MKOH solution. Calculate the pH after the following additions of the KOH solution: (a) 0.0 mL, (b) 5.0 mL, (c) 10.0 mL, (d) 12.5 mL, (e) 15.0 mL. (25 points)

            (Reference: Chang 16.35)

  1. Referring to the following table:

Indicator

Color in Acid

Color in Base

pH Range*

Thymol blue

Red

Yellow

1.2–2.8

Bromophenol blue

Yellow

Bluish purple

3.0–4.6

Methyl orange

Orange

Yellow

3.1–4.4

Methyl red

Red

Yellow

4.2–6.3

Chlorophenol blue

Yellow

Red

4.8–6.4

Bromothymol blue

Yellow

Blue

6.0–7.6

Cresol red

Yellow

Red

7.2–8.8

Phenolphthalein

Colorless

Reddish pink

8.3–10.0

*The pH range is defined as the range over which the indicator changes from the acid color to the base color.

Specify which indicator or indicators you would use for the following titrations:  (9 points)

  1. HCOOH versus NaOH
  2. HCl versus KOH
  3. HNO3versus CH3NH2

(Reference: Chang 16.43)

  1. The pKaof butyric acid (HBut) is 4.7. Calculate Kbfor the butyrate ion (But). (5 points)

(Reference
Answer #1

& Finst questions : Determination ut ut i - Here the given acid is acetic acid (CHg. CooH). It is a weak acid. For any weak aNow given that I the concentration c = 0.4 M And. pka of CH2 LOUH = 4.76 (data collected). Now from equation ③ we have - pH =[salt] pH = pka + log [ Acid the get of Now, Here given, TAcid) = CH2, Cro H = 0-4 M [salt] = [CH3COONa] = 0.2 m for acetic a(Reference: Change 16-5) A buffen solution is the mixture of solutions of weak acid and its salt. Now HU and Hq so both are sfor these type of Solution - [salt] рон ерко тоГоле) Kb = ionisation constant of weak base Again, pH = 14 -POH Now given - [BPlease,,,,,for the other answers post them as another questions.....